3 A motor has 5 speeds varying from X rpm to Y rpm in a Geometric Progression.                                                                  

Calculate the common ratio

Make a table of the speeds.


1
Expert's answer
2021-10-12T06:07:28-0400

Initial Speed = x

Final Speed = y

Say the geometric ratio = r

In case of geometric progression, the above case be represented as


"x, xr, xr^2, xr^3, xr^4, , xr^5, y"


Therefore, it can be said that

"x r ^5 = y"

Therefore,

"r^5 = \frac{y}{x}"

Or, "r = (\frac{y}{x})^{1/5}"

Therefore, the common ratio is given by: "r = (\frac{y}{x})^{1/5}"